# PropMolFlow: Property-guided Molecule Generation with Geometry-Complete Flow Matching

### v3

@@ Methods
We compare PropMolFlow against several recent property-guided molecular generation methods based on diffusion models and flow matching, including EEGSDE, GCDM, GeoLDM, JODO, and EquiFM [27, 25, 8, 44, 21]. In this work, we focus exclusively on conditional generation results; unconditional generation performance is reported in earlier studies. Language-model-based approaches are excluded from this comparison, as state-of-the-art models in that category rely on fundamentally different molecule representations and training data regimes. Our comparison includes both structural-validity and property-specific evaluation metrics. For EEGSDE, we benchmark PropMolFlow against the canonical conditional model with a scaling factor of 1 ($s=1$[27]. Although EEGSDE provides DFT-based validation for five out of six properties, it does so on only 100 generated molecules, which limits the statistical significance of those results.
EEGSDE reported the atomic and molecular stability results for its conditional generation, JODO reported molecular stability, and GCDM reported PoseBusters validity for the conditional generation of GCDM and GeoLDM models—no other baseline conditional models provided these metrics. Because EEGSDE and GCDM released their model checkpoints, we sampled from those and computed all structural validity metrics on the resulting structures; for GeoLDM, we used the checkpoints bundled with GCDM. EquiFM checkpoints were unavailable, so we trained its conditional models ourselves following information in [21] (training details in Supplementary Section LABEL:sec:more_details).
Except for the closed-shell ratio, all structural validity metrics depend on bond-order information, which baseline models (other than JODO) do not output. We therefore assigned bond orders using the distance-based cutoffs introduced by Hoogeboom [7]. Although recent work [23] suggests that the bond orders can probably be optimized through the OpenBabel program [45]. We did not apply such refinements to the baseline samples.
For property metrics, PropMolFlow is directly compared to prior methods on the ID tasks. OOD results are not directly comparable due to a lack of baseline generative models and property prediction models testing this regime, but we validate our results with DFT calculations to assess physical fidelity.
@@
Hyperparameters, Training and Inference Details
PropMolFlow uses the model architecture detailed in Supplemental Section LABEL:sec:appendix_model. PropMolFlow models on the QM9 dataset were trained with 8 molecule update blocks. Atoms contain 256 hidden scalar features and 16 hidden vector features. Edges contain 128 hidden features. QM9 PropMolFlow models were trained with 2000 epochs. Training and inference for PropMolFlow models and all baseline models used a single NVIDIA A100-SXM4 graphic card with 80GB of memory, with a batch size of 128. PropMolFlow models can be trained in about 4–5 days, and EquiFM models were trained with 2500 epochs, which took around 3.5 days.
During inference, PropMolFlow models use Euler’s method with 100 evenly spaced time steps to integrate the learned velocity field. The number of atoms for each generated molecule is first sampled from the QM9 atom count distribution, and the property values are then sampled according to the number of atoms to ensure consistency with the dataset’s statistics, as used in previous property-conditioning works [7, 8, 25].
GVP regressor details.
To be self-consistent, we trained property regressors using Graph Neural Networks based on GVPs. GVP property predictors are trained on a disjoint dataset versus that used for training PropMolFlow models, hence avoiding data leakage between the generative models and property prediction models. We appended an MLP layer that takes the final node scalar features as input to predict the target property. The parameters of GVP regressors are optimized by minimizing a mean squared error loss function. Separate models were trained for each of the six tested properties.
Acknowledgment
@@
Data and code availability
Code and data of this work can be found on github: https://github.com/Liu-Group-UF/PropMolFlow.

### v4

@@ Methods
In this work, we focus exclusively on conditional generation results; unconditional generation performance is reported in earlier studies. Language-model-based approaches are excluded from this comparison, as state-of-the-art models in that category rely on fundamentally different molecule representations and training data regimes. For EEGSDE, we benchmark PropMolFlow against the canonical conditional model with a scaling factor of 1 (s=1s=1[bao2023]. Although EEGSDE provides DFT-based validation for five out of six properties, it does so on only 100 generated molecules, which limits the statistical significance of those results. Because EEGSDE, JODO and GCDM released their model checkpoints, we sampled molecules from those and computed all structural validity metrics on the resulting structures; for GeoLDM, we used the checkpoints bundled with GCDM. EquiFM checkpoints were unavailable, so we trained its conditional models by ourselves following information in [song2023] (training details in Supplementary Section LABEL:sec:more_details).
Three intuitive baselines are included for comparison in the ID task. The “Random (Upper-bound)” intuitive baseline fully shuffles property values, thereby removing any correlation between molecular structures and properties, and serves as an upper bound on error. The “# Atoms” baseline uses only the number of atoms as a predictor for molecular properties, capturing coarse size-dependent trends. It is implemented as a simple neural network with one hidden layer [hoogeboom2022]. The “QM9 (Lower-bound)” baseline corresponds to a pretrained EGNN regressor provided by Hoogeboom et al. [hoogeboom2022], trained directly on QM9 and serving as a lower bound on achievable error. Improvement over the “Random” baseline indicates that the conditional generation effectively integrates property information. Surpassing the “# Atoms” baseline suggests that the generative model captures structural features beyond simple atom count when generating new molecules.
Evaluation metrics
Structural validity metrics.
Our comparison includes both structural-validity, inference-efficiency and property-specific evaluation metrics. Structural validity metrics include atomic stability, molecule stability, RDKit validity, ‘uniqueness and validity’, PoseBusters validity, and closed-shell ratio. Atom stability is given by the proportion of atoms with correct valency. To account for atomic charges, an atom is considered to have the correct valency if its formal charge balances its explicit valency, as defined in RDKit [rdkit]. For instance, a nitrogen atom should carry a +1 formal charge if it has a valency of 4. Molecule stability defines the proportion of molecules in which all atoms are stable and the molecule is charge neutral (zero net formal charges). RDKit validity refers to the ratios of molecules that pass RDKit’s sanitization check. ‘Uniqueness and validity’ is defined as the proportion of molecules that are RDKit valid and unique in their SMILES representation. PoseBusters validity corresponds to the fraction of molecules that pass many de novo chemical and structural validation tests, such as all atom connectivity, valid bond lengths and angles, and absence of internal clashes [buttenschoen2024]. We used the first 10 columns of the PoseBusters outputs in CSV formats as the criteria for results shown in Fig. 2 and Supplementary Table LABEL:tab:standard_metrics. A molecule is closed-shell valid if it has an even number of valence electrons. Since all molecules in QM9 are closed-shell, any generated open-shell species are considered invalid. Except for the closed-shell ratio, all structural validity metrics depend on bond-order information, which baseline models (other than JODO) do not output. We therefore assigned bond orders using the distance-based, optimized cutoffs introduced by Hoogeboom [hoogeboom2022]. Although recent work [vignac2023] suggests that the bond orders can probably be optimized through the OpenBabel program [O'Boyle2011], we did not apply such refinements to the baseline samples. To assess the chemical similarity between generated molecules and a reference training set, we employed a fingerprint-based approach using RDKit. For each molecule, a Morgan fingerprint (circular fingerprint) was computed with a radius of 2 and a 2048-bit representation. Pairwise similarity between each generated molecule and all training molecules was computed using the Tanimoto similarity metric on their respective Morgan fingerprints. For each generated molecule, the maximum similarity score and the corresponding most similar training molecule were recorded. Lastly, inference efficiency is given by the wall-clock time required to generate 10,000 molecules.
Property metrics. PropMolFlow is directly compared to prior methods on the ID tasks. In this task, a conditional generative model was first used to generate molecules conditioned jointly on property values and atom counts sampled from the same distribution as the training data (for example, QM9). These sampled property values and atom counts served as the input conditions. A separate property predictor/regressor was then applied to estimate the properties of the generated molecules. In PropMolFlow, we used an SE(3) GVP network for molecule generation and, correspondingly, trained a separate GVP regressor for each of the six molecular properties. 10000 molecules are generated to evaluated the performance of conditional generation. OOD results are not directly comparable to baseline models, but we validate our results with DFT calculations to assess physical fidelity.
RMSD and property sensitivity.
Given the same chemical compositions and element-type ordering, the RMSD between any two molecules (M0 and M1) is the minimum distance between these two structures considering the translational and rotational symmetry operations, which can be expressed as:
Where SE(3) represents the group of operations that respect the translational and rotational symmetry. To estimate the property sensitivity to structural relaxation, we selected the structures whose RMSDs are larger than 0.03 Å—RMSDs lower than this threshold are considered very close to the relaxed structures, and in Eq. (19), we define the normalized property sensitivity $\chi_{\rm{DFT}}$’ to allow comparisons between different properties.
where $\delta q_{\rm{DFT}}$ is the change of DFT-calculated property upon relaxation, and the normalizer ($q_{0.99}-q_{0.01}$) is the difference between the 99th quantile and the 1th quantile of the property distribution of QM9 training data. This range is chosen over the min-max range because it is robust to outliers.
@@
GVP regressor details
To be self-consistent, we trained property regressors using Graph Neural Networks based on GVPs. GVP property predictors are trained on a disjoint dataset versus that used for training PropMolFlow models, hence avoiding data leakage between the generative models and property prediction models. We appended an MLP layer that takes the final node scalar features as input to predict the target property. The parameters of GVP regressors are optimized by minimizing a mean squared error loss function. Separate models were trained for each of the six molecular properties.
DFT
All DFT calculations were performed using the Gaussian 16 package (Gaussian 16, Revision C.01) [g16]. The B3LYP hybrid functional [becke1993, lyp1988] together with the 6-31G(2df,p) basis set was employed for single-point calculations, and geometric optimizations. Single-point calculations were used unless otherwise specified. Structural relaxations were performed for the property $C_{v}$ in most cases, except in the interpolation study. Relaxations were performed to evaluate changes in both properties and geometries, as shown in Fig. 3. DFT calculations that failed to converge were discarded. We released the DFT-evaluated molecules in both the ID and OOD tasks with and without structural optimization at the Zenodo repository [zenodo_propmolflow].
Molecule filtering
We selected 1000 molecules out of 10000 samples structures in the ID task or use the directly generated 1000 structures in the OOD task for DFT evaluations. multi-step filtering is essential to ensure that the structures are chemically valid and suitable for downstream applications. We filtered the 1000 molecules by five criteria: Molecule stability, RDKit validity, PoseBusters validity, closed-shell validity, and multi-fragment check. The multi-fragment check filters out molecules that with more than one disconnected fragment. For molecules passing all filters, property values are evaluated using both a GVP property predictor and DFT calculations. We also filtered out molecules whose DFT calculations did not converge, resulting in no less than 90% of and 78–94% of the molecules remaining for the subsequent analyses in the respective ID and OOD tasks (Supplementary Table LABEL:tab:no_sp_dft).
Hyperparameters and training details
PropMolFlow models on the QM9 dataset were trained with 8 molecule update blocks. Atoms contain 256 hidden scalar features and 16 hidden vector features. Edges contain 128 hidden features. QM9 PropMolFlow models were trained with 2000 epochs. For each property embedding method, 6 model checkpoints with lowest validation losses were saved and the top-3 model checkpoints were deposited to the Zenodo repository [zenodo_propmolflow]. Inference for PropMolFlow models and all baseline models used a single NVIDIA A100-SXM4 graphic card with 80GB of memory, with a batch size of 128. Training of PropMolFlow models use both Nvidia A100-SXM4 and 2080ti graphic cards. PropMolFlow models can be trained in 2–3 days with A100 and 4–5 days with 2080Ti cards, and EquiFM models were trained with 2500 epochs on A100, which took around 3.5 days.
Data availability
All data can be found in the Zenodo repository [zenodo_propmolflow], including revised QM9 SDF data, SDF files for generated raw molecules, DFT-calculated structures and their molecular properties saved in the extxyz format, full property MAE and structural validity results in CSV format for all saved PropMolFlow checkpoint models, sampled structures and their full PoseBusters results for all baseline models, and EGNN property-predictor pretrained models and notebook examples to use the property predictors.
Code availability
Our PropMolFlow implementation is available at https://github.com/Liu-Group-UF/PropMolFlow and on Zenodo [propmolflow_code].
Acknowledgments